OpenStudy (anonymous):

Find x–y if x/y=1/7. i don't understand why you would do x+y=1? but that seems crucial in solving the problem..

OpenStudy (anonymous):

sorry i still don't see the full solution?

OpenStudy (anonymous):

sigh the answer key says that its -3/4

OpenStudy (anonymous):

i appreciate your help though.

OpenStudy (anonymous):

this is their explanation on the attached file. it makes no sense to me but maybe it will make sense to you?

OpenStudy (anonymous):

basically you just cross multiply and simplify...

OpenStudy (anonymous):

yeah we end up with -6x but the answer is -3/4

OpenStudy (anonymous):

did you check my attachment?

OpenStudy (anonymous):

Yes, I did. Cross multiplying x/y=1/7 gives y = 7x. If you substitute 7x for y in the equationx + y = 1, you then get x + 7x = 1, 8x = 1. AND x=1/8

OpenStudy (anonymous):

THEN......

OpenStudy (anonymous):

where did the x+y come from?

OpenStudy (anonymous):

thats the main issue

OpenStudy (anonymous):

well.. x+y=1..

OpenStudy (anonymous):

okay, I see your problem. Did your teacher go over this at all with you?? You may want to talk to them about this because its kind of confusing.... What you have to do is switch the signs around because then it isn't the "stem" equation. THEN one you find the answer to that part of the equation, you can go ahead and use the x-y part of the equation

OpenStudy (anonymous):

I can't really explain it. Its one of those things that you have to show full out....do you get what I mean?

OpenStudy (anonymous):

uh, so do you mind drawing it?

OpenStudy (anonymous):

@Hero can you please take a look at this question you may be able to explain better than me.

OpenStudy (anonymous):

ok don't feel pressured..this problem sucks i know.

OpenStudy (anonymous):

Haha, I get you....sorry, 'tis cool...

OpenStudy (anonymous):

Umm, do you mind waiting until tomorrow for a drawing? My mom is kinda pissed Im not asleep...you know how they get when they are in their "moods"...lol. sorry about that Youngin...have you tried Googleing your question?

OpenStudy (anonymous):

don't worry about it. get to bed before your mom kills you

OpenStudy (anonymous):

Okay, thanks. I bookmarked this page so I can do it before I go to school tomorrow for you, Okay?

OpenStudy (anonymous):

lol its cool. don't worry too much.

OpenStudy (anonymous):

okay, thanks. But that drawing will be here sooner or later. PROMISE lol, i like helping people out

OpenStudy (anonymous):

@Harkirat look at this question and see if you can give Youngin a good/better explanation than I did

OpenStudy (anonymous):

k, now I am going to bed, lol

OpenStudy (anonymous):

ok -just adding my 2 cents. basically when you initially asked this question you did not say that prior information supplied to you was that x + y = 1. That is crucial information

OpenStudy (anonymous):

we all agree, 7x =y once we do cross multiplication. now rewrite the given; x + y = 1 to x + 7x =1 8x = 1 x =1/8 with that substitute back into x + y = 1 1/8 + y = 1 y = 1- 1/8 y = 8/8-1/8 y = 7/8 now you are looking for x -y 1/8 - 7/8 = -6/8 = -3/4  let me know if this is clear now.